A 0.10 kg model rocket's engine is designed to deliver an impulse of 6.0 N*s. If the rocket engine burns for 0.75 s, what is the average force does the engine produce?


1
Expert's answer
2021-06-14T14:58:46-0400
"J=F_{avg}\Delta t,""F_{avg}=\dfrac{J}{\Delta t}=\dfrac{6.0\ N\cdot s}{0.75\ s}=8.0\ N."